### v1

# Reasoning is All You Need for Video Generalization: A Counterfactual Benchmark with Sub-question Evaluation

Abstract
Counterfactual reasoning is crucial for robust video understanding but remains underexplored in existing multimodal benchmarks. In this paper, we introduce COVER (COunterfactual VidEo Reasoning), a multidimensional multimodal benchmark that systematically evaluates MLLMs across the abstract-concrete and perception-cognition dimensions. Beyond prior multimodal benchmarks, COVER decomposes complex queries into structured sub-questions, enabling fine-grained reasoning analysis. Experiments on commercial and open-source models reveal a strong correlation between sub-question accuracy and counterfactual reasoning performance, highlighting the role of structured inference in video understanding. Furthermore, our results suggest a key insight: enhancing the reasoning capability of models is essential for improving the robustness of video understanding. COVER establishes a new standard for assessing MLLMs’ logical reasoning abilities in dynamic environments.
Reasoning is All You Need for Video Generalization: A Counterfactual Benchmark with Sub-question Evaluation
Qiji Zhou11footnotemark: 1 1, YiFan Gong11footnotemark: 1 2, Guangsheng Bao1, Hongjie Qiu2, Jinqiang Li2 Xiangrong Zhu2, Huajian Zhang1, Yue Zhang22footnotemark: 2 1 1School of Engineering, Westlake University 2College of Computer Science and Technology, Hangzhou Dianzi University {zhouqiji, baoguangsheng, zhanghuajian, zhangyue}@westlake.edu.cn {gongyifan, qiuhongjie, lijinqiang, zhuxiangrong}@hdu.edu.cn
1 Introduction
In recent years, the rapid advancement of large language models (LLMs) has spurred growing interest in multimodal large language models (MLLMs) Hurst et al. (2024); Anthropic (2024); Chen et al. (2025); Zhang et al. (2024, 2025); Wang et al. (2024); Wu et al. (2024b). Various early benchmarks have been proposed to assess multimodal understanding ability of MLLMs, particularly in static images Fu et al. (2024a); Hudson and Manning (2019); Liu et al. (2024); Yu et al. (2024). More recently, benchmarks involving complex images and dynamic videos have emerged to evaluate MLLM’s capabilities in temporal reasoning, spatio-temporal recognition, and object detection Fu et al. (2024b); Li et al. (2024b, 2023). Despite these advances, these benchmarks often overlook counterfactual reasoning, which is a critical component for evaluating inference in complex and realistic environments. As a result, they fall short of providing a comprehensive assessment of MLLMs’ reasoning capabilities.
Counterfactual reasoning, which posits hypothetical alternatives to observed realities, is pivotal for advanced video inference and is closely tied to out-of-distribution generalization (Yang et al., 2023; Bao et al., 2025). Previous work has attempted to construct counterfactual queries for images and videos (Li et al., 2024d, e, c; Patel et al., 2022; Wu et al., 2023). Most existing multimodal counterfactual benchmarks tend to focus on assessing subtask-specific robustness of reasoning ability (Li et al., 2024e; Wu et al., 2024c, 2023). However, they do not assess the underlying factors that contribute to the robustness of these reasoning capabilities. Such benchmarks often lack a systematic progression from abstract to concrete dimensions and from low-level perception to high-level cognition, limiting their ability to comprehensively capture multimodal reasoning processes in MLLMs. Furthermore, these benchmarks rarely investigate how robust video understanding interacts with stepwise reasoning in dynamic environments, leaving a gap in our assessment of advanced inference skills.
To bridge this gap, we propose COVER, a counterfactual video reasoning benchmark driven by a multidimensional abstraction level evaluation mechanism. Unlike existing multimodal counterfactual benchmarks, which often focus on multitask-oriented questions, COVER systematically classifies tasks into four quadrants. We define specific tasks for each quadrant to evaluate MLLMs’ diverse reasoning capabilities in complex video scenarios. Beyond merely posing counterfactual questions, COVER introduces a sub-question reasoning mechanism derived from necessary conditions, enabling a deeper evaluation of performance across MLLMs. This approach allows us to establish a connection between the accuracy of intermediate steps and the overall robustness of counterfactual reasoning. As shown in Figure 1, when asked to determine whether a boy completes a series of actions in a specified order, COVER decomposes the problem into multiple steps, each representing a necessary condition. For instance, sub-question Q1 may inquire about which action occurs first in the reversed video, while sub-question Q2 targets the final action. This structured approach not only helps evaluate how a model adapts to event-sequence changes but also reveals its strengths and weaknesses in extracting and synthesizing critical information under counterfactual assumptions. By encompassing a broad range of abstraction levels, COVER stands as the most comprehensive dataset of its kind, paving the way for more rigorous and holistic evaluations of MLLMs’ dynamic and counterfactual reasoning capabilities.
Building on the COVER benchmark, we conducted a series of systematic experiments using both open-source and commercial closed-source models of varying scales. Our results indicate a strong positive correlation between the models’ sub-question accuracy and performance in counterfactual reasoning and robust video understanding. The findings underscore the tight linkage between sophisticated inference capabilities and high-level video comprehension. Furthermore, we examine how automatically generated versus human-guided sub-question decomposition (chain-of-thought, CoT Wei et al. (2022)) influences complex reasoning and identifies the key factors impacting inference accuracy in MLLMs. Through these experiments, COVER offers valuable insights into how structured reasoning can enhance the robustness of video understanding by constructing a sub-question–based counterfactual video QA benchmark across multiple levels of abstraction and thoroughly evaluating mainstream MLLMs’ logical reasoning abilities.
| Benchmark | Video | Q&A | Qs Source | CF | SQP | PCD | ACD |
| CoFCA (Wu et al., 2024a) | ✗ | ✓ | H&A | ✓ | ✓ | ✗ | ✗ |
| CFMM (Li et al., 2024e) | ✗ | ✓ | H | ✓ | ✗ | ✗ | ✗ |
| Video-MME (Fu et al., 2024b) | ✓ | ✓ | H | ✗ | ✗ | ✓ | ✗ |
| CRIPP-VQA (Patel et al., 2022) | ✓ | ✓ | H | ✓ | ✗ | ✗ | ✗ |
| VITATECS (Li et al., 2024c) | ✓ | ✗ | H&A | ✓ | ✗ | ✗ | ✗ |
| COVER (ours) | ✓ | ✓ | H&A | ✓ | ✓ | ✓ | ✓ |
2 Related Work
MLLMs and Their Evaluation. Recent advancements in Multimodal Large Language Models (MLLMs) have significantly enhanced their ability to process multimodal data, such as text and images, or text and video. Specialized evaluation benchmarks have been developed to assess models’ understanding of static images, such as MM-Vet Yu et al. (2024), MME Fu et al. (2024a), MMBench Liu et al. (2024), and GQA Hudson and Manning (2019), which primarily focus on visual comprehension. Extending these to dynamic multimodal reasoning, video-based benchmarks like Video-MME Fu et al. (2024b), Mvbench Li et al. (2024b), and SEED-Bench Li et al. (2023) evaluate models’ temporal and contextual understanding. These benchmarks play a crucial role in evaluating the capabilities of MLLMs in dynamic and evolving environments, highlighting the growing emphasis on multimodal reasoning across diverse tasks.
Chain-of-Thought and Counterfactual Reasoning in MLLMs. Chain-of-Thought (CoT) reasoning Wei et al. (2022) decomposes complex problems into sequential steps, improving logical reasoning. In MLLMs, multimodal CoT variants Zhang et al. (2023b); Zheng et al. (2023) have shown notable improvements in reasoning across modalities. Furthermore, counterfactual reasoning, which investigates how modifying specific conditions affects outcomes, has gained increasing attention. Previous studies have evaluated LLMs on text-based counterfactual tasks Wu et al. (2024c, a) and visual counterfactual Q&A Li et al. (2024e). ACQUIRED Wu et al. (2023) offers a comprehensive taxonomy for counterfactual scenarios, enriching this area of research. Additionally, CoFCA Wu et al. (2024a) incorporates sub-problem evaluation to assess evidence integration and multi-step reasoning capabilities. CoT and counterfactual reasoning provide a more systematic and effective approach to multimodal reasoning, significantly enhancing the models’ ability to handle complex reasoning tasks.
Generalization of Multimodal Reasoning and Video Counterfactual Tasks. Despite significant contributions from video counterfactual benchmarks like CRIPP-VQA, which focuses on physical attributes, VITATECS, which emphasizes captioning, and ACQUIRED’s limited counterfactual taxonomy Li et al. (2024c); Patel et al. (2022), their scope remains limited.
To address these gaps, COVER introduces sub-question decomposition, allowing the evaluation of perceptual and cognitive reasoning in counterfactual video tasks. As shown in Table 1, COVER spans four dimensions—distinguishing between abstract and concrete object attributes and perceptual and cognitive demands offering a more comprehensive evaluation framework than existing benchmarks.
3 The COVER Benchmark
This section provides a comprehensive overview of the construction of COVER. We introduce its data partitioning framework designed to evaluate MLLM reasoning ability across four complementary dimensions. Next, we describe the data curation process, which domain experts have rigorously validated to ensure the high quality and reliability of the benchmark.
3.1 Benchmark Definition
As illustrated in Figure 2, we categorize the 13 benchmark tasks into four quadrants based on the abstract-concrete and perceptual-cognitive dimensions. Abstract-Perception: (1) Emotion: Understanding and recognizing emotional states. Concrete-Perception: (2) Counting: Quantity recognition and calculation. (3) Color: Perceiving object colors. (4) Direction: Sensing motion trends. (5) Size: Identifying object dimensions. (6) Shape: Perceiving object shapes. (7) Material: Recognizing object materials. (8) Location: Detecting object positions. Concrete-Cognition: (9) Action Recognition: Identifying specific actions. (10) Object Recognition: Recognizing specific objects. Abstract-Cognition: (11) Action Prediction: Forecasting future actions. (12) Procedure Understanding: Comprehending sequential processes and logic. (13) Social Relation: Understanding social relationships.
Division of Abstract and Concrete Scenes. This distinction reflects a functional differentiation within cognitive representation systems. Neuroscientific studies Katja Wiemer-Hastings and Xu (2005) suggest that concrete concepts rely heavily on multi-modal perceptual simulations (e.g., object shape, material), while abstract concepts are primarily represented through language-mediated symbolic operations. Abstract tasks often require integrating non-perceptual information, such as contextual encoding for emotion recognition or constructing temporal causal models for action prediction.
Division of Perception and Cognition. Perception involves the initial reception of external stimuli through sensory organs, converting them into neural signals that provide raw environmental data. Cognition, built upon perception, refers to the further processing of these signals, encompassing higher-level mental functions such as memory, attention, language comprehension, problem-solving, and reasoning. This distinction underscores different stages of information processing, with perception forming the foundation upon which cognitive functions are built.
3.2 Data Construction
To construct COVER, we carefully selected a diverse range of open-source and research-available video sources, including but not limited to Sigurdsson et al. (2016); Yi et al. (2020); Xie et al. (2024); Tan et al. (2020); Shahroudy et al. (2016); Patraucean et al. (2024); Zhang et al. (2023a); Gao et al. (2017); Jang et al. (2017); Wang et al. (2019); Krantz et al. (2020). These sources encompass various real-world scenarios, ranging from daily activity recognition to complex scene understanding. As shown in Appendix Figure 7, we collected 146 videos and designed 150 aspect-specific QA pairs, each of which underwent dual-team review for validation. To ensure balanced coverage across the four quadrants, we expanded the seed data using GPT-generated instances (720-760 per quadrant) to mitigate any potential biases. The detailed statistical findings are comprehensively presented in Figure 3. The frame count of videos in COVER ranges from 16 to 1739, with an average of 294.34 frames. We finally constructed 2,923 high-quality counterfactual question-answer pairs. Each question-answer pair consists of an original question, which presents no hypothetical context, and a counterfactual question, which introduces situational assumptions and sub-questions that enable granular reasoning analysis.
Eight annotators further validated the dataset and checked logical consistency to ensure the reasoning relied solely on the video content. Additionally, three experts cross-valid the dataset (see Appendix Table 6) to confirm the structural balance.
| GPT-4o | 70.26 | 45.93 | 56.94 |
| GPT-4o-mini | 67.32 | 51.47 | 55.94 |
| Claude-3.5-Sonnet | 63.60 | 38.04 | 49.40 |
| Gemini-1.5-Pro | 74.82 | 49.64 | 63.76 |
| Gemini-1.5-Flash | 73.90 | 48.75 | 62.52 |
| Gemini-2.0-Flash | 77.18 | 46.90 | 62.92 |
| InternVL2.5-78B | 76.74 | 59.46 | 67.23 |
| LlaVA-Video-72B | 64.35 | 56.04 | 61.54 |
| InternVL2.5-26B | 75.40 | 51.08 | 62.65 |
| InternVL2.5-8B | 74.31 | 57.75 | 61.65 |
| VideoLlama3-8B | 73.04 | 51.25 | 60.09 |
| LlaVa-OV-7B | 62.74 | 51.80 | 56.42 |
| LLaVA-Video-7B | 60.52 | 51.93 | 55.11 |
| Qwen2-VL-7B | 71.83 | 46.90 | 58.40 |
| VILA-U-7B | 60.01 | 38.42 | 47.32 |
| VILA1.5-7B | 60.25 | 57.34 | 53.18 |
4 Experiments
In this section, we conduct a thorough evaluation of MLLMs of varying scales on the COVER dataset. We assess model performance across four key quadrants, ranging from perception to cognition and from abstract to concrete. Then, we evaluate their capabilities in counterfactual reasoning, direct reasoning, and sub-question-guided reasoning across multiple subtasks. The results include a detailed comparison of open-source and commercial closed-source models of different sizes.
4.1 Settings
To ensure a comprehensive and in-depth evaluation, we selected a diverse set of representative MLLMs, including both commercial closed-source models such as GPT-4o Hurst et al. (2024), Claude Anthropic (2024), and Gemini Team et al. (2024), as well as prominent open-source models like InternVL2.5 Chen et al. (2025), LLAVA-Video Zhang et al. (2024), LLaVA-OV Li et al. (2024a), Qwen2-VL Wang et al. (2024), VideoLLaMA3 Zhang et al. (2025), vila-u Wu et al. (2024b), and VILA Lin et al. (2023). These models represent a broad spectrum of parameter scales and technical characteristics, providing a well-rounded reflection of technological development in multimodal learning. We use , , and to evaluate the original question accuracy, counterfactual question accuracy, and sub-question accuracy of MLLM in video understanding, using the average score from at least three runs. We adopted a consistent frame extraction strategy for the video inputs, sampling 16 frames per segment, ensuring that all models are evaluated under identical conditions. In addition, we will explore the effect of different sampling strategies in Chapter 5.
Models AC (%) CC (%) CP (%) AP (%) GPT-4o 71.05 41.81 41.70 74.87 43.65 68.36 69.95 42.62 50.52 65.01 55.65 63.97 GPT-4o-mini 62.29 52.40 42.97 76.32 54.37 66.49 64.62 40.85 44.78 65.56 58.26 65.96 Claude-3.5-sonnet 56.92 37.01 35.55 70.11 42.33 61.77 60.03 32.88 40.08 66.94 39.81 56.81 Gemini 1.5 Pro 69.49 44.49 53.36 82.14 51.98 72.78 71.76 43.93 56.81 75.48 57.99 69.54 Gemini 1.5 Flash 70.48 45.34 52.23 82.01 49.34 71.51 70.67 42.02 51.90 72.04 58.26 71.36 Gemini 2.0 Flash 74.29 44.36 51.38 83.99 47.75 72.84 74.22 38.74 58.26 75.90 57.71 66.84 InternVL2.5-78B 72.88 59.60 57.67 80.95 63.62 75.62 75.99 58.25 63.65 76.86 56.20 70.07 LLaVA-Video-72B 53.11 54.94 53.14 65.34 60.45 67.03 67.94 52.39 53.49 70.66 56.20 70.01 InternVL2.5-26B 71.05 47.74 50.53 80.95 58.99 72.17 76.13 47.20 60.12 73.14 50.00 65.61 InternVL2.5-8B 69.77 58.62 49.96 80.95 64.55 71.02 73.94 55.80 54.66 72.18 51.79 68.19 VideoLLama3-8B 68.08 45.62 49.68 81.35 54.89 68.36 72.99 50.75 51.62 69.28 53.44 67.90 LLaVA-ov-7B 54.66 51.69 47.49 62.96 53.04 61.77 64.53 49.66 49.48 68.60 52.75 64.73 LLaVA-Video-7B 50.14 55.23 44.52 61.64 50.53 60.50 63.57 52.52 49.97 66.39 49.59 63.03 Qwen2-VL-7B 65.96 49.15 48.41 82.14 43.39 67.03 71.21 45.57 50.52 67.49 49.72 65.02 VILA-U-7B 58.19 39.83 38.87 63.10 41.93 54.51 59.07 37.93 37.94 59.50 33.88 55.34 VILA1.5-7B 54.80 55.93 39.29 66.93 62.30 63.52 55.25 58.53 44.64 63.64 52.34 61.91
4.2 Main Results
As shown in Table 2, Gemini-2.0-Flash ( 77.18%) and InternVL2.5-78B ( 76.74%) rank as the top two models, demonstrating their strong foundational video understanding. The lower scores of VILA-U-7B ( 60.01%) and LLaVA-Video-7B ( 60.52%) highlight the limitations of smaller models in processing long sequences. InternVL2.5-78B ( 59.46%) shows significant dominance in handling conditional reasoning and complex contexts. Notably, counterfactual questions cause sharp accuracy drops compared to the original questions: GPT-4o (-24.33%) and Gemini-1.5-Pro (-25.18%), indicating that most models struggle with counterfactual reasoning.
Most models exhibit higher than (e.g., Claude-3.5-Sonnet 49.40% vs. 38.04%, LLaVA-Video-72B 61.54% vs. 56.04%). This suggests better stability in localized reasoning tasks than in holistic tasks, where error accumulation impacts performance. In the Appendix, we provide detailed case demonstrations in Figure 8.
Closed-source Model Performance. As shown in Table 3, Gemini 1.5 Pro demonstrates strong dominance in both concrete cognition ( 82.14%) and abstract perception tasks ( 75.48%). Gemini 2.0 Flash excels in abstract perception ( 75.90%) and concrete perception tasks ( 74.22%), showcasing strong capabilities in handling high-complexity perceptual tasks.
Open-source Model Performance. As shown in Table 3, InternVL2.5-78B leads in abstract cognition ( 72.88%) and concrete perception tasks ( 58.25%), reflecting a deep understanding of abstract concepts and complex logic. Lightweight models like Qwen2-VL-7B perform well in concrete cognition ( 82.14%) but face limitations in abstract tasks ( 65.96% in AC) due to their smaller parameter size, revealing distinct capabilities across model types. Commercial models, such as the Gemini series, maintain strong performance in concrete cognition and abstract perception tasks but generally fall behind open-source models in counterfactual reasoning. Most models struggle with counterfactual reasoning, with only InternVL2.5-7BB and VILA1.5-7B showing some task-specific advantages, highlighting the need for targeted optimization in conditional hypothesis modeling.
4.3 Sub-question Guideline
| Model | Without CoT | With CoT | Guide-CoT | |
| GPT-4o-mini | 51.47 | 58.62 | 57.93 | 68.07 |
| InternVL2.5-78B | 59.46 | 60.42 | 58.33 | 68.29 |
| LlaVA-Video-72B | 56.04 | 56.24 | 53.51 | 63.12 |
| InternVL2.5-8B | 57.75 | 57.06 | 52.41 | 57.75 |
| VideoLlama3-8B | 51.25 | 52.82 | 53.06 | 52.79 |
| LLaVA-Video-7B | 51.93 | 51.42 | 51.39 | 54.12 |
| Qwen2-VL-7B | 46.90 | 50.36 | 45.71 | 50.88 |
We propose Guide-CoT to study the influence of different reasoning paths on model performance through human-annotated sub-problems. We design comparative experiments between CoT and Guide-CoT to analyze how automatically generated sub-questions from CoT versus manually annotated sub-questions affect model reasoning capabilities.
Comparing the Without CoT and CoT approaches based on Table 4, we find that the of most models under CoT significantly exceeds the Without CoT baseline, such as Qwen2-VL-7B (+3.46%) and GPT-4o-mini (+7.15%), which indicates that CoT enhances reasoning processes, particularly in more complex tasks.
However, examining Guide-CoT results reveals that manually designed sub-questions may not always lead to substantial improvement over automatically generated ones, as seen with GPT-4o-mini’s of 57.93% under Guide-CoT, slightly lower than the 58.62% under CoT. This does not imply the ineffectiveness of manual sub-questions but suggests that model behaviors may not always align with human-designed reasoning paths, potentially due to task complexity or the nature of the sub-questions themselves. We hypothesize that manually provided sub-questions could introduce extraneous patterns or "pseudo-features" that are not directly relevant to the reasoning task, leading to a subtle reduction in performance.
The column in Guide-CoT represents sub-questions containing standard answers. For InternVL2.5-78B, reaches 68.29%, showing an 8.63% improvement over the no-CoT baseline, compared to CoT’s mere 0.96% enhancement (59.46% → 60.42%). This demonstrates that providing complete answers significantly enhances reasoning accuracy, especially in complex or multi-step tasks. Sub-questions with standard answers help the model better integrate information and validate reasoning steps, leading to improved consistency and performance. In Appendix Figure 9, we provide detailed case demonstrations to further explore these observations and examine the relationships between reasoning paths and task complexities.
The results from our experiments strongly support the notion that reasoning plays a pivotal role in model robustness and generalization. Our study extends these insights by demonstrating that multimodal models, especially in the context of video tasks, rely heavily on robust reasoning capabilities for effective generalization. The significant performance improvements observed with counterfactual reasoning and sub-question decomposition highlight that models’ ability to handle complex, conditional, and dynamic contexts is crucial for their robustness, a finding not fully explored in prior research.
Frames InternVL2.5-1B InternVL2.5-2B InternVL2.5-4B InternVL2.5-8B 2 66.16 35.61 55.27 65.31 44.20 54.99 72.56 48.31 60.88 71.26 58.50 60.07 4 68.32 34.72 55.52 68.83 42.11 58.84 74.41 46.49 61.79 73.35 58.47 60.96 8 68.94 35.10 55.11 68.22 41.43 55.75 75.03 45.60 61.79 74.14 57.06 61.60 16 69.76 35.89 55.19 70.07 40.68 55.49 75.61 45.23 61.63 74.31 57.75 61.65 32 69.04 36.50 55.04 70.13 39.69 55.48 75.54 45.09 60.96 74.10 57.03 61.42 64 68.18 37.39 54.80 68.90 40.06 55.44 74.41 46.56 60.70 74.20 58.09 61.30
5 Analysis
In this chapter, we begin by analyzing the impact of video frame sampling rates on MLLMs’ video understanding and reasoning abilities. We then proceed with an in-depth examination of MLLMs’ robustness and logical reasoning performance.
5.1 Ablation Study of Video Frames
As shown in Table 5, as the parameter size of LLMs increases, there is a rising trend in , , and . For instance, with 16 frames, the InternVL2.5-1B model achieves , , and of 69.76%, 35.89%, and 55.19% respectively. The InternVL2.5-2B scores 70.07%, 40.62%, and 55.49%, while the InternVL2.5-4B reaches 75.61%, 45.23%, and 61.68%, indicating that larger LLMs have enhanced capabilities in handling complex problems. Under the same vision tower settings, shows a clear upward trend as the number of frames increases. For example, the InternVL2.5-8B’s rises from 71.26% at 2 frames to 74.20% at 64 frames. However, tends to decrease with more frames. The InternVL2.5-2B’s drops from 44.20% at 2 frames to 40.06% at 64 frames. Models with more parameters generally perform better in , , and , highlighting the significant role of LLMs in multimodal reasoning. Additionally, increasing visual information (by raising the frame count) can enhance , but excessive visual information, especially in complex or counterfactual reasoning scenarios, may impair the model’s reasoning ability, leading to a decline in .
5.2 An Analysis of Robustness and Logical Reasoning Capability in MLLMs
The ability of MLLMs to answer original questions serves as a key indicator of their overall understanding capabilities, while performance on sub-questions reveals single-step reasoning proficiency. A notable observation is the strong Pearson correlation between and reaches 0.836, indicating a strong connection between model understanding and reasoning capabilities. Furthermore, as shown in Table 2, the correlation between and is 0.608. These moderately strong correlations indicate that a model’s ability to comprehend original questions plays a fundamental role in enabling effective step-by-step reasoning. similarly, the correlation between and suggests that models with a higher understanding capability tend to perform better when solving decomposed sub-questions, reinforcing the notion that comprehension and reasoning are interdependent. However, the moderate correlation between and suggests that counterfactual reasoning involves additional complexities, making it a more challenging task than single-step reasoning.
Heatmaps in Figure 4 reveal distinct performance patterns across the different quadrants, and the relationship between comprehension, step-by-step reasoning, and counterfactual inference. In the abstract cognition quadrant, which includes high-level reasoning tasks, the performance drop from to is minimal, and the transition to remains stable. This suggests that while sub-question reasoning is important, models still maintain strong performance for abstract reasoning tasks even when counterfactual assumptions are introduced. Conversely, in the concrete perception quadrant, a more pronounced performance decline is observed from to , and even further to . This indicates that sub-question reasoning has a more significant impact on counterfactual reasoning in perception-based tasks, where models struggle to decompose complex perceptual information into manageable steps.
Overall, these findings suggest that counterfactual reasoning is inherently more complex than single-step reasoning, particularly in perception-intensive tasks where models must infer causal relationships rather than relying on pattern recognition alone. The relatively stable transition between and in abstract cognition tasks implies that models leverage their existing conceptual understanding to handle counterfactual scenarios more effectively. Improving counterfactual reasoning in perception-heavy tasks remains a key challenge, likely requiring enhancements in reasoning mechanisms and causal inference capabilities.
6 Conclusion
This paper introduces COVER, a comprehensive benchmark for counterfactual video reasoning, evaluating MLLMs across four dimensions: abstract-concrete and perception-cognition. By decomposing complex queries into structured sub-questions, COVER enables a detailed analysis of reasoning performance, highlighting the link between sub-question accuracy and overall counterfactual reasoning. Our findings emphasize that enhanced reasoning capabilities are essential for improving model performance in dynamic, real-world video tasks. COVER sets a new standard for evaluating logical reasoning in multimodal video tasks, driving the development of models that can effectively reason across dynamic video scenarios.
Limitations
COVER offers a novel benchmark for counterfactual video reasoning, but some limitations exist. First, while it focuses on video reasoning, its applicability to other multimodal tasks, such as image or text reasoning, remains unexplored. Second, COVER relies on sub-question decomposition, and automated methods may not always match human-designed questions, especially in complex scenarios. Finally, while we demonstrate COVER’s effectiveness on various models, further validation across different model architectures and real-world tasks is needed to assess its generalizability.
Ethical Considerations
COVER is designed with ethical considerations in mind, aiming to enhance counterfactual reasoning in video understanding while ensuring fairness, transparency, and responsible AI development. We acknowledge the ongoing challenges in bias mitigation, fairness, and environmental sustainability and encourage the broader research community to collaborate in addressing these concerns. By establishing COVER as an open and structured evaluation benchmark, we aim to promote robust and ethical AI advancements in multimodal reasoning.
We ensured that the human annotators were compensated with fair remuneration, which exceeded the local minimum wage standards, reflecting the value of their work. Furthermore, we took steps to ensure that the annotation process did not pose any risks to their physical or mental well-being. The tasks were designed to be manageable, and we provided adequate support to ensure a safe and respectful working environment.
In this study, AI was used solely for data augmentation and grammar/typo correction, with no involvement in generative or creative tasks. We carefully considered potential risks to ensure AI usage did not compromise the originality or transparency of the research.
Appendix A Appendix
A.1 Data Construction Details
In this section, we present additional details on COVER construction, including information about the annotation agreements, data augmentation prompts, and process flow.
| Aspect | A | B | C | Average |
| Data Quality | 4 | 4 | 5 | 4.3 |
| Data Diversity | 5 | 4 | 5 | 4.7 |
| Relevance | 4 | 5 | 4 | 4.3 |
| Annotation Quality | 4 | 5 | 5 | 4.7 |
| Dataset Usability | 4 | 4 | 4 | 4 |
| Innovation | 5 | 5 | 4 | 4.7 |
Three annotators were recruited to evaluate COVER across multiple dimensions, with the resultant assessments systematically compiled in Table 6, ensuring comprehensive evaluation coverage. The methodological workflow for data augmentation is schematically outlined in Figure 7.
The schematic framework outlined in Figure 5 delineates the methodology employed for contextual data augmentation, leveraging the generative capabilities of GPT-4oHurst et al. (2024) to construct domain-specific instructional prompts.
A.2 Additional Results
In this section, we present additional experiments on COVER. The comprehensive evaluation framework delineated in Table 8 presents granular performance metrics across 13 meticulously defined tasks.
GPT-4o exhibited notable vulnerability in the Procedure Understanding task. While it attained a respectable raw accuracy of 78.17%, its counterfactual accuracy plummeted to 28.97%, representing a precipitous decline of 49.2%. This substantial drop suggests that the performance of GPT-4o in understanding procedures may be overly reliant on surface-level features. Counterfactual perturbations, such as changes in conditions, can severely disrupt its reasoning capabilities, thereby highlighting a robustness limitation of the model when handling complex tasks.
Figure 6 (a) depicts the relationship between and across different models, with a purple regression line characterizing the functional correlation between mean and mean . Figure 6 (b) demonstrates the association between and across different models, with a red regression line characterizing the functional correlation between mean and mean . The bivariate correlation analysis delineated in Figure 6 demonstrates statistically significant covariation patterns (r = 0.836) between semantic comprehension and multi-step reasoning capabilities in MLLMs.
A.3 Sample Reaults on Test Time Long Reasoning Models
As illustrated in Figure 10, the reasoning model QVQ-72B-Preview Team (2024), equipped with a built-in Chain-of-Thought (CoT) mechanism, exhibits human-aligned reasoning patterns. Its cognitive process integrates detailed scenario descriptions, systematic elimination of implausible options (e.g., excluding candidates A/B/C), and rigorous conclusion verification. In contrast, InternVL2.5-78B employs a CoT mechanism that presents answers in a bullet-point format without explanatory justification, reflecting weaker anthropomorphic reasoning characteristics.
However, the discrepancy in Table 7 (QVQ-72B-Preview: 59.33% < InternVL2.5-78B: 71.33%) suggests that contemporary reasoning models may rely more on memorization than on structured reasoning. InternVL2.5-78B’s concise response paradigm appears to leverage rapid pattern recognition and information retrieval, leading to superior accuracy. While QVQ-72B-Preview’s elaborate reasoning workflow better approximates human cognition, potential redundancies or logical inconsistencies may reduce answer precision.
Table 7 further indicates that InternVL2.5-78B achieves a substantial lead in the metric (70.80%), significantly outperforming QVQ-72B-Preview (58.76%) and Claude-3.7-sonnect (46.72%). This performance hierarchy remains consistent across models when evaluated on the metric: InternVL2.5-78B (70.00%) > QVQ-72B-Preview (69.33%) > Claude-3.7-sonnect (46.00%). Empirical evidence suggests a statistically significant positive correlation between reasoning capability () and comprehension ability (). In addition, under the CoT paradigm, reasoning capability demonstrates stronger generalization, exhibiting a positive correlation with performance on human-annotated essential logical sub-problems, thereby reinforcing the intrinsic relationship between logical reasoning and generalizability.
Moreover, the reasoning processes of models such as QVQ frequently generate sub-problem content that aligns with human-annotated data, which to some extent suggests that the inferential patterns of test-time long-reasoning models demonstrate closer correspondence with human cognitive intuition. For instance, in the Figure 11 the analytical content regarding the opening and closing scenes of videos (highlighted in blue font) exhibits precise alignment with the manually curated sub-problems in the upper-right annotation (specifically addressing inquiries about video commencement and conclusion scenarios), thereby empirically validating this cognitive congruence.
| QVQ-72B-Preview | 69.33 | 59.33 | 58.76 |
| Claude-3.7-sonnect | 46.00 | 59.33 | 46.72 |
| InternVL2.5-78B | 70.00 | 71.33 | 70.80 |
| VILA1.5-13B | 65.33 | 44.67 | 53.65 |
Model Type Task Action Procedure Social Action Object Color Counting Direction Location Material Shape Size Emotion Prediction Understanding Relation Recognition Recognition GPT-4o 65.20 78.17 69.00 74.87 74.87 92.23 75.25 50.88 70.59 79.12 72.00 52.88 65.01 41.41 28.97 56.33 44.65 42.67 37.86 40.59 33.33 42.86 59.34 58.00 29.81 55.65 51.85 22.82 52.43 69.54 67.09 51.94 47.52 56.90 48.96 58.08 55.56 36.08 63.97 GPT-4o-mini 50.22 72.22 63.32 78.61 74.08 84.47 70.30 52.63 57.98 71.43 62.00 56.73 65.56 44.05 51.19 62.01 56.42 52.36 26.21 39.60 36.84 59.66 52.75 53.00 17.31 58.26 53.16 19.44 58.85 68.03 64.82 38.35 38.12 53.88 47.72 53.89 53.44 28.85 65.96 Claude-3.5-Sonnet 43.61 63.10 63.32 66.31 73.82 79.61 68.32 48.25 52.10 63.74 66.34 45.19 66.94 39.21 33.33 38.86 43.85 40.84 36.89 19.80 35.96 40.34 37.36 40.59 18.27 39.81 46.19 15.87 46.68 62.54 60.93 39.81 24.26 48.28 48.55 46.11 37.70 34.13 56.81 Gemini-1.5-Pro 54.63 80.95 71.62 83.42 80.89 84.47 73.27 61.40 76.47 81.32 67.33 59.62 75.48 46.70 29.76 58.52 45.45 58.38 46.60 37.62 35.96 42.86 54.95 55.45 36.54 57.99 57.52 39.68 64.38 72.58 72.99 73.30 43.56 61.21 58.09 59.88 55.50 45.67 69.54 Gemini-1.5-Flash 53.74 85.32 70.74 82.62 81.41 82.52 70.30 57.02 70.59 79.12 69.31 68.27 72.04 45.81 34.92 56.33 49.20 49.48 41.75 37.62 33.33 41.18 64.84 53.47 25.96 58.26 61.87 32.94 63.94 73.28 69.60 46.60 43.07 62.93 54.77 62.87 55.50 37.98 71.36 Gemini-2.0-Flash 60.35 86.90 74.24 87.97 80.10 90.29 69.31 64.04 78.99 81.32 70.30 66.35 75.90 42.29 36.51 51.97 44.12 51.31 20.39 39.60 31.58 37.82 57.14 56.44 31.73 57.71 60.78 35.52 59.51 73.16 72.49 69.90 59.41 65.95 53.53 58.08 58.64 42.31 66.84 InternVL2.5-78B 67.84 75.00 75.55 79.68 82.20 94.17 82.18 52.63 76.47 76.92 83.17 69.23 76.86 43.61 76.19 57.21 65.51 61.78 87.38 37.62 47.37 75.63 61.54 57.43 39.43 56.20 62.09 44.64 67.70 76.90 62.28 79.13 69.80 66.38 58.09 62.28 59.69 50.48 70.07 LLaVA-Video-72B 43.17 50.79 65.50 60.70 69.90 85.44 69.31 51.75 73.11 74.73 61.39 61.54 70.66 44.93 59.92 59.39 63.10 57.85 62.14 42.57 47.37 66.39 53.85 51.49 41.35 56.20 59.26 32.94 69.47 67.56 66.46 63.59 52.97 61.21 45.23 55.69 53.40 43.27 70.01 InternVL2.5-26B 57.27 78.58 76.42 82.35 79.58 91.26 74.26 62.28 85.71 74.73 78.22 66.35 73.14 47.14 45.24 51.09 60.43 57.59 59.23 25.74 45.61 60.50 57.14 25.00 25.00 50.00 59.91 61.08 62.39 71.18 73.24 65.05 56.44 68.97 58.09 61/08 50.96 50.96 65.61 InternVL2.5-8B 55.51 75.00 78.17 81.28 80.63 90.29 70.30 63.16 78.99 74.63 74.26 66.35 72.18 48.02 76.19 49.78 71.39 57.85 84.47 36.63 53.51 70.59 59.34 55.45 28.85 51.79 55.99 29.37 66.81 69.89 72.24 52.43 52.97 60.34 53.53 56.89 54.97 51.44 68.19 VideoLLama3-8B 52.42 81.75 68.56 80.48 82.20 94.17 70.30 63.16 81.51 70.33 68.32 62.50 69.28 35.68 53.97 46.29 55.08 54.71 66.02 42.57 42.11 64.71 58.24 48.51 32.69 53.44 49.45 33.93 67.48 67.91 68.84 57.77 39.60 58.62 49.79 53.89 54.45 47.12 67.90 LLaVA-ov-7B 48.90 48,81 66.81 60.43 65.45 86.41 63.37 44.74 63.03 72.53 61.39 63.46 68.60 43.61 64.29 45.85 55.35 50.79 59.22 42.57 45.61 52.94 60.44 57.43 30.77 52.75 50.11 30.36 63.94 62.78 60.68 54.85 42.08 53.45 45.64 50.90 48.69 50.96 64.73 LLaVA-Video-7B 50.66 35.71 65.50 56.15 67.02 83.50 67.33 41.23 58.82 74.72 64.36 59.62 66.39 44.93 73.02 45.85 59.09 42.15 73.79 42.57 48.25 60.50 58.24 48.51 35.58 49.59 48.58 29.96 56.64 62.19 58.67 55.34 41.09 56.03 43.15 54.49 52.88 48.08 63.03 Qwen2-VL-7B 44.49 84.12 67.25 84.22 80.10 88.35 70.29 57.89 73.94 74.73 65.34 69.23 67.49 42.29 58.73 45.41 44.92 41.88 56.31 21.78 42.11 58.82 60.44 45.54 33.65 49.72 53.37 30.16 63.72 67.33 66.71 43.20 42.08 59.48 51.87 52.69 51.83 51.44 65.02 VILA-U-7B 45.37 73.02 54.59 66.31 59.95 81.55 61.39 54.39 71.43 47.25 47.52 47.11 59.50 22.47 53.17 42.36 44.39 39.53 45.63 45.54 23.68 43.70 35.16 35.64 36.54 33.88 38.78 34.33 44.03 52.16 57.04 41.26 22.28 39.22 40.66 35.93 42.93 42.31 55.34 VILA1.5-7B 52.86 50.40 61.57 67.65 66.23 61.17 56.44 40.35 40.34 71.43 60.40 62.50 63.64 28.64 85.71 50.22 68.45 56.29 86.41 57.43 49.12 76.47 51.65 41.58 44.23 52.34 34.86 24.80 59.96 61.73 65.45 48.06 31.19 50.86 42.74 47.31 45.55 46.63 61.91
A.4 Examples of Sub-question Guidelines
Figure 8 illustrates how sub-question errors propagate to counterfactual question failures. In Figure 9, we observe that subtle errors in the reasoning process lead to reasoning failures, highlighting the model’s sensitivity to the integrity of its reasoning steps.